Find the domain of r(t) and the value of r(t0).1, t 2; t0 = 1


1
Expert's answer
2021-09-01T09:21:04-0400
"\vec r(t)=\langle t^2,\dfrac{1}{t}, \sqrt{1-t}\rangle"

"t\not=0, 1-t\geq0"

"Domain: (\infin, 0)\cup(0, 1]"


"t_0=1"


"\vec r(t_0)=\vec r(1)=\langle 1^2,\dfrac{1}{1}, \sqrt{1-1}\rangle"

"=\langle 1,1, 0\rangle"

"\vec r(t_0)=\vec r(1)=\vec i+\vec j"